OpenStudy (anonymous):

PLEASE HELP WITH SOLVING EQUATION WITH RATIONAL EXPRESSION !!!!CONSIDER THE FOLLOWING EQUATION X+7/X+2=-1/4

OpenStudy (mathmale):

x + 7/x + 2 = -1/4 is to be solved for x (i. e., we want to find the value of x, if any, that satisfies this equation). Let's start by eliminating fractions. One fraction in this equation is 7/x, with denominator x.. The other fraction is -1/4. Thus, the lowest common denominator is 4x. Multiply all terms of the equation by 4x. This results in: 4x^2 +28 + 8x = -x Move -x to the left side of the equation: 4x^2 + 9x + 28 = 0. Unfortunately, this quadratic equation does not have real roots. My conclusion would be that there is no real x value that satisfies the equation you've provided. Would you mind checking whether or not you've copied the problem statement correctly?